For the following equation, state the value of the discriminant and then describe the nature of the solutions.
-8x^2-9x+19=0
What is the value of the discriminant?
Which one of the statements below is correct?
a. The equation has two real solutions.
b. The equation has one real solution.
c. the equation has two imaginary solutions.

Answer by stanbon(75887)

-8x^2-9x+19=0
---
a = -8 ; b = -9 ; c = 19
------
What is the value of the discriminant?
b^2-4ac = 81 - 4*-8*19 > 0
-------
Which one of the statements below is correct?
Since the discriminant is positive, "a" is correct.
-----------------------
Cheers,
Stan H.
-----------------------
a. The equation has two real solutions.
b. The equation has one real solution.
c. the equation has two imaginary solutions.